for the function f(x)=x^2+3+2, what is f(x+h)=


1
Expert's answer
2021-10-17T16:44:14-0400

For the function "f(x)=x^2+3x+2", let us find "f(x+h):"


"f(x+h)=(x+h)^2+3(x+h)+2=x^2+2hx+h^2+3x+3h+2\\=x^2+ (3+2h)x+h^2+3h+2."